HUBBLE SPACE TELESCOPE DAILY REPORT #1032 PERIOD COVERED: 0800 EST Daily Status Report as of 348/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Complete S/C 4761 (SMOV FOS Recovery to Hold) The FOS was recovered from Safe mode with no problems seen in the telemetry. The last pressure reading on the vacuum gage was 0.16E-02 mT. No anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: No Guide Star acquisitions were done during this period. 2.2 FHST Updates: No FHST updates were done during this period. 2.3 Significant HSTARs: None 2.4 Operations Note: Transition from the Health and Safety SMS to SMS-347 was successfully completed at 347/2100Z and the first STScI SMS based loads have begun execution as planned. The NSSC-1 pointer was set, and the GHRS aliveness test was completed. The LGBU was successfully uplinked at 348/0854Z. An HGBU will be uplinked after the FHST Calibrations are complete. HST Servicing Mission Activities continue as planned and all subsystems remain normal. Normal performance continues for all the PCS hardware and software. The newly replaced gyros and magnetometers continue to show normal behavior. All six rate gyros are powered "on". FHST/RGA Scale Factor alignment data collection during six vehicle slews with gyros in high mode has been completed. The solar arrays were successfully slewed for the first time since RMS release. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete Three Sets of FHST1/2/3/4 5253 (Internal FHST/RGA Scale Factor Alignment) 3.2 Complete Six Sets of S/C 4762 (FOS Pressure Measurement) 3.3 Complete S/C 4756 (GHRS SMOV Restore Door Configuration) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Planned SMOV phase activities. /DMH -END-